# Chaotic stochastic resonance in Mackey–Glass equations

### v3

@@ Abstract
Stochastic resonance (SR) is observed as switching dynamics between two quasi-stationary states in stochastic Mackey–Glass equations. We identify a new form of SR, chaotic SR, characterized by positive Lyapunov exponents, arising from the coexistence of SR and stochastic chaos. Unlike stable SR, which exhibits negative Lyapunov exponents, the resonance point for chaotic SR precedes the zero-crossing point of the largest Lyapunov exponent. Furthermore, we provide a theoretical estimate of the resonant periods for both stable and chaotic SR based on a linear mode analysis around an unstable fixed point.
Stochastic resonance (SR) is a noise-induced phenomenon in which the response of a nonlinear system to an external force is enhanced by the addition of an optimal level of noise. SR can also arise in time-delayed systems, where the delayed feedback plays the role of an external force. In this study, we investigate the stochastic Mackey–Glass equation, a chaotic time-delayed system subject to noise, and identify two distinct types of SR, stable SR and chaotic SR. Since SR has often been examined from the perspective of stochastic processes, the underlying dynamical properties remain unclear. To clarify these properties, we apply random dynamical system theory to characterize the dynamical structure of the stochastic system. Stable SR is characterized by random point attractors, whereas chaotic SR is characterized by random strange attractors.
@@ I Introduction
Adding noise to a deterministic dynamical system can induce qualitative changes in its behavior, known collectively as noise-induced phenomena. Stochastic resonance (SR) Gammaitoni et al. (1998) is a representative example, in which the system’s response to an external force is enhanced by an optimal level of noise. SR was originally investigated in bistable systems subjected to noise and periodic forcingBenzi, Sutera, and Vulpiani (1981); BENZI et al. (1982), where the addition of noise was found to enhance the periodicity of the system’s dynamics. Since then, SR has been observed in a broad range of nonlinear systems. Even without periodic forcing, noise-induced periodicity can emerge ( i.e., coherence resonance Pikovsky and Kurths (1997)). Likewise, the interplay between time delay and noise can generate a resonance in the absence of external forcing, known as delayed SR Ohira and Sato (1999); Tsimring and Pikovsky (2001); Masoller (2002). We refer to these phenomena collectively as SR in a broad sense. Based on the idea that deterministic chaos may act similarly to noise, related SR-like behavior in chaotic systems without external noise has also been studied, commonly referred to as chaotic resonance Carroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Krawiecki et al. (2001); Jüngling et al. (2008).
SR has also been reported in chaotic systems under both external noise and periodic forcing Anishchenko, Neiman, and Safanova (1993); Anishchenko, Safonova, and Chua (1992). However, these studies have largely relied on the framework of stochastic processes, leaving the dynamical nature of resonance in chaotic regimes insufficiently clarified. To address this gap, we employ random dynamical system theory Arnold (1998); Chekroun, Simonnet, and Ghil (2011), which enables analysis of the intrinsic dynamical structure of stochastic systems. Within this framework, we identify stable SR characterized by random point attractors, and chaotic SR characterized by random strange attractors.
This paper is organized as follows. Section II.1 introduces the stochastic Mackey–Glass equation and its phenomenology. Section III presents numerical and analytical results for both stable and chaotic SR. Section IV provides concluding remarks.
@@ II.1 Stochastic Mackey–Glass equations
where $\tau>0$ denotes the delay time, $a>0$ is the strength of delayed feedback, $b>0$ is the decay rate, and $c$ is the shape parameter of the delayed feedback. The initial condition is defined as $x(t)=\phi(t)$ for $t\in[-\tau,0)$ and is assumed to be constant unless otherwise specified. The MG can exhibit multiple positive Lyapunov exponents, and the effective dimension of the dynamics increases as $\tau$ grows Doyne Farmer (1982).
@@
where $\sigma$ represents the noise intensity and $W_{t}$ denotes a Wiener process. The parameters are fixed at the standard values $b=0.1$ and $c=10$, while $a$ and $\sigma$ serve as control parameters. The delay time is chosen to be sufficiently larger than the response time Mensour and Longtin (1998) $1/b$ and is set to $\tau=90$ unless stated otherwise.
@@ II.2 Bifurcation in the stochastic Mackey–Glass equation
We briefly summarize the phenomenology of the dynamics in the MG and SMG. In the bifurcation diagram of the MG (Fig. 1 (a)), the Poincaré section, $dx/dt=0$, of the attractor in Eq. (1) is plotted as a function of the delayed feedback strength $a$ (blue dots). In the bifurcation diagram of the SMG (Fig. 1 (b)), projections of the random pullback attractors are shown as black dots (see Appendices A and B for details on pullback attractors and bifurcation diagram). The first and second Lyapunov exponents of the MG and SMG as functions of $a$ at $\sigma=0.0$ and $\sigma=0.15$ are shown in Fig. 1 (c) (blue and black, respectively). For the MG, the origin is stable for $0<a<0.1$. A pitchfork bifurcation occurs at $a=0.1$, giving rise to two symmetric fixed points with respect to the origin. As $a$ increases, a Hopf bifurcation emerges at $a\simeq 0.125$, producing a limit cycle. A bifurcation to chaos follows at $a\simeq 0.138$, with high-dimensional chaos appearing around $a\simeq 0.144$. In the SMG, the zero-crossing point of the first Lyapunov exponent occurs at $a=a_{c}\simeq 0.175$, beyond which stochastic chaosChekroun, Simonnet, and Ghil (2011) is observed. At $a=a_{d}\simeq 0.195$, the second Lyapunov exponent becomes positive, indicating the emergence of high-dimensional stochastic chaos.
In summary, although the MG becomes multi-stable via a pitchfork bifurcation, the SMG exhibits a single attractor with two quasi-stable states as two deterministic attractors are connected by external noise. As $a$ increases, the Lyapunov exponents of the SMG grow monotonically, leading to stochastic chaos and later to high-dimensional stochastic chaos. Notably, these stochastic bifurcations occur at significantly larger $a$ values than in the deterministic system.
Figure 1: (a) Bifurcation diagram of the MG with $\sigma=0$. (b) Bifurcation diagram of random pullback attractors of the SMG with $\sigma=0.15$. (c) First and second Lyapunov exponents for $\sigma=0$ (blue points) and $\sigma=0.15$ (black points). In the bifurcation diagram, random pullback attractors are computed using $1.0\times 10^{2}$ initial conditions with a pullback time $t_{p}=4\times 10^{4}$. We designate $a_{1}=0.14$ as a representative example of a weakly nonlinear regime and $a_{2}=0.30$ as that of a strongly nonlinear regime. At $a=a_{1},a_{2}$, low- and high-dimensional deterministic chaos are observed, respectively. The values, $a_{c}$ and $a_{d}$, indicate the zero-crossing points of the first and second Lyapunov exponents in the presence of noise. Numerical simulations are conducted using the Euler–Maruyama method with a time step of $\Delta t=0.1$, integrating for $10^{8}$ steps.
@@ II.3 Stochastic resonance in the stochastic Mackey–Glass equation
Beyond the deterministic pitchfork bifurcation point ($a>0.1$), SR appears as switching dynamics between two quasi-stable states, associated with point, periodic, and strange attractors in the MG. Figure 2 illustrates SR for a regime with deterministic periodic attractors. Without noise, two symmetric limit cycles exist with respect to the origin (Fig. 2 (a)), each with a period of approximately $2\tau$. In the presence of noise, switching between these limit cycles emerges, and periodicity is enhanced at an optimal noise intensity. Figure 2 (c) shows the power spectra with and without noise (black and blue lines, respectively). Under optimal noise intensity, the peaks at $f=f_{n}^{*}\simeq n/\tau$ $(n=1,2,\ldots)$ reveal clear harmonics. The power of the primary resonant frequency $f_{1}^{*}$ reaches a maximum at the optimal noise intensity $\sigma=\sigma^{*}$, confirming the presence of SR. We refer to $\sigma^{*}$ as the resonance point (Fig. 2 (d)).
Figure 2: (a) Two deterministic attractors projected onto $(x(t),x(t-\tau))$ for $a=0.126$. The inset shows a magnified view of the limit cycle. (b) Time series for the deterministic case, $\sigma=0$ (blue dots), and the resonance case, $\sigma=\sigma^{*}$ (black dots). (c) Power spectra for $\sigma=0$ (blue lines) and the optimal noise intensity $\sigma=\sigma^{*}$ (black lines). (d) Power of the primary resonant frequency $f_{1}^{*}$ as a function of $\sigma$.
@@ III.1 Stable and chaotic SR
To characterize the dynamical properties of SR, we compute the Lyapunov exponents of the SMG (see Appendix A for the definition of Lyapunov exponent in random dynamical systems). We examine two representative cases: one where the deterministic attractors exhibit low-dimensional chaos ($a=a_{1}=0.14$), and one where they exhibit high-dimensional chaos ($a=a_{2}=0.30$). In both cases, switching dynamics between two quasi-stable states appears, and the power spectrum peaks at $f^{*}_{n}\simeq n/\tau$ $(n=1,2,\ldots)$ are enhanced at the resonance point (Fig. 3). The sign of the largest Lyapunov exponent at the resonance point distinguishes the two types of SR: a negative exponent corresponds to stable SR (Fig. 3 (b)), whereas a positive exponent corresponds to chaotic SR (Fig. 3 (d)).
Figure 3: Power spectra for $\sigma=0$ (blue lines) and $\sigma=\sigma^{*}$ (black lines) at (a) $a=a_{1}$ and (c) $a=a_{2}$. Power of the primary resonant frequency $f_{1}^{*}$ (black points) as a function of $\sigma$ at (b) $a=a_{1}$ and (d) $a=a_{2}$. The largest Lyapunov exponent at $a=a_{1}$ and the first six Lyapunov exponents at $a=a_{2}$ are also shown in (b) and (d), respectively (red points).
In the case of stable SR, increasing the noise intensity causes $\lambda_{1}$ to become negative at $\sigma=\sigma_{0}$, at which point the chaotic attractor collapses into a random point attractor. As $\sigma$ is increased further, $\lambda_{1}$ attains a local maximum near the resonance point. Because the zero-crossing point $\sigma_{0}$ precedes the resonance point $\sigma^{*}$, chaos vanishes at resonance, rendering the system effectively analogous to the classical SR model with a double-well potential and periodic forcing Benzi, Sutera, and Vulpiani (1981). For sufficiently large $\sigma$, $\lambda_{1}$ decreases again, as trajectories stay almost always in contracting regions outside the attractors. In the case of chaotic SR, $\lambda_{1}$ initially increases and peaks near the resonance point. Here, $\sigma^{*}$ precedes $\sigma_{0}$, indicating that SR occurs while the system still possesses a positive Lyapunov exponent and thus retains stochastic chaos. Chaotic SR can be regarded as a class of noise-induced phenomena that enhances the characteristic periods of stochastic chaos. The ordering of $\sigma^{*}$ and $\sigma_{0}$ provide a clear diagnostic: $\sigma_{*}>\sigma_{0}$ for stable SR and $\sigma_{*}<\sigma_{0}$ for chaotic SR.
@@ III.2 Visualization of stable and chaotic SR
Previous studies on SR in chaotic systems have generally not distinguished chaotic SR from stable SR Anishchenko, Neiman, and Safanova (1993); Anishchenko, Safonova, and Chua (1992); Carroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Krawiecki et al. (2001); Jüngling et al. (2008), largely due to the lack of a framework rooted in random dynamical system theory. The distinction becomes clear when visualizing the dynamics using space–time representations Arecchi et al. (1992); Yanchuk and Giacomelli (2017) and random pullback attractors (see Appendix A for the construction of random pullback attractors). A space–time representation maps the time series into a two-dimensional plane by introducing a discrete step $n\in\mathbb{Z}$ and a memory space $s\in[0,\tau)$. Both stable SR and chaotic SR appear as traveling waves in the memory space $s$ (Fig. 4, left column). Notably, chaotic SR exhibits pronounced periodicity, with chaotic structures clearly visible inside the traveling waves, indicating the coexistence of SR and stochastic chaos. When the largest Lyapunov exponent is negative, the random pullback attractor collapses to a random point attractor because initially close trajectories converge and eventually synchronize. Denoting a distribution of all possible trajectories at time $t$ as $\rho_{t}$, stable SR can be interpreted as pseudo-periodic motion of $\rho_{t}$ on the random point attractor (Fig.4 (b)). Conversely, when Lyapunov exponents are positive, the random pullback attractor becomes a random strange attractor. In this setting, chaotic SR corresponds to pseudo-periodic motion of $\rho_{t}$ on a random strange attractor (Fig. 4 (d)). This pseudo-periodic motion of $\rho_{t}$ is related to the problem of statistical periodicity Lasota and Mackey (1987); Losson and Mackey (1995). The precise relationship between chaotic SR and statistical periodicity remains an open question and presents an interesting direction for future study.
@@
Figure 4: Space–time representation at the resonance point for (a) stable SR ($a=a_{1},\sigma=0.162$) and (c) chaotic SR ($a=a_{2},\sigma=0.104$). Both cases display traveling waves in memory space $s\in[0,\tau)$. The angle $\theta$ denotes the propagation direction of the traveling wave relative to the orthogonal direction, defined by $\tan\theta=\epsilon$. Using the measured angles $\theta_{1}$ (stable SR) and $\theta_{2}$ (chaotic SR), we obtain $\theta_{2}/\theta_{1}\simeq 0.786$, indicating that wave propagation is slower in the chaotic SR regime. (b) and (d) show snapshots of the random pullback attractor at the resonance point computed from $2.0\times 10^{4}$ initial conditions after a pullback time $t_{p}=2\times 10^{4}$, projected onto $(x(t),x(t-\tau))$.

### v4

@@ Abstract
Stochastic resonance (SR) manifests as switching dynamics between two quasi-stationary states in the stochastic Mackey–Glass equation. We identify chaotic SR, arising from the coexistence of resonance and chaos in stochastic dynamics. In contrast to classical SR, which is described by a random point attractor with a negative largest Lyapunov exponent, chaotic SR is described by a random strange attractor with a positive largest Lyapunov exponent. We observe chaotic SR in the Mackey-Glass equation as well as chaotic SR in the Duffing equation and the underdamped FitzHugh-Nagumo equation, demonstrating the universality of this phenomenon across a broad class of strongly nonlinear random dynamical systems.
Stochastic resonance (SR) is a noise-induced phenomenon in which the response of a nonlinear system to an external forcing is enhanced at an optimal noise intensity. In the present work, we collectively refer to these phenomena as “SR,” in the broad sense of resonance phenomena arising in stochastic dynamics. SR can also arise in time-delayed systems, where the delayed feedback plays the role of an external forcing. In this study, we investigate the stochastic Mackey–Glass equation, a chaotic time-delayed dynamics subject to noise, and identify two distinct types of SR, namely stable SR and chaotic SR. To clarify these properties, we employ concepts in random dynamical systems theory, which enables us to directly analyze the intrinsic dynamical structure in stochastic dynamics. Stable SR is described by random point attractors with a negative largest random LE, whereas chaotic SR is described by random strange attractors with a positive largest random LE. We observe chaotic SR in the Mackey-Glass equation as well as chaotic SR in the Duffing equation and the underdamped FitzHugh-Nagumo equation, demonstrating the universality of this phenomenon across a broad class of strongly nonlinear random dynamical systems.
@@ I Introduction
Adding noise to a deterministic dynamical system can induce qualitative changes in its behavior, known collectively as noise-induced phenomena. Stochastic resonance (SR) Gammaitoni et al. (1998) is a representative example, in which the response of a nonlinear system to an external forcing is enhanced at an optimal noise intensity. SR was originally investigated in bistable systems subjected to noise and periodic forcingBenzi, Sutera, and Vulpiani (1981); BENZI et al. (1982), where the addition of noise enhances the periodicity of the system’s dynamics. Since then, SR has been observed in a wide variety of nonlinear systems. Even in the absence of periodic forcing, noise-induced periodicity can emerge, a phenomenon known as coherence resonance Pikovsky and Kurths (1997). The interplay between time delay and noise can also generate a coherence resonance in time-delayed systems Ohira and Sato (1999); Tsimring and Pikovsky (2001); Masoller (2002). In the present work, we collectively refer to these phenomena as “SR,” in the broad sense of resonance phenomena arising in stochastic dynamics.
SR has also been reported in chaotic systems Carroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Jüngling et al. (2008); Anishchenko, Neiman, and Safanova (1993); Anishchenko, Safonova, and Chua (1992); Zakharova et al. (2010). Most previous studies have been conducted within the framework of deterministic dynamical systems theory, since the stochastic process approachesGammaitoni et al. (1998); Benzi, Sutera, and Vulpiani (1981); McNamara and Wiesenfeld (1989); Wiesenfeld et al. (1994); Gingl, Kiss, and Moss (1995) does not focus on the characterization of the underlying dynamical structure of the system. Deterministic dynamical systems approaches primarily address resonance phenomena in the absence of external noise by interpreting deterministic chaos as an effective noise sourceCarroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Jüngling et al. (2008). These studies have shown that SR can be described in terms of crisis. Furthermore, extensions of stochastic approaches to deterministic crises and noise-induced crisesSommerer, Ott, and Grebogi (1991) have demonstrated that chaotic systems can exhibit stochastic multiresonanceVilar and Rubi (1997), in which the response to periodic forcing is maximized at multiple distinct noise intensities Krawiecki et al. (2001); Matyjaśkiewicz et al. (2003).
In this paper, we employ random dynamical systems theory, which enables a direct analysis of the intrinsic dynamical structure in stochastic systems, including random attractors, stochastic bifurcations, and stability measured by random Lyapunov exponents (LEs). Within this framework, we identify stable SR, described by random point attractors with a negative largest random LE, and chaotic SR, described by random strange attractors with a positive largest random LE, in Mackey–Glass equations.
This paper is organized as follows. Section II introduces the stochastic Mackey–Glass equation and its phenomenology. Section III presents numerical and analytical results on stable and chaotic SR. Section  IV provides concluding remarks.
@@ II.1 Stochastic Mackey–Glass equations
where $\tau>0$ denotes the delay time, $a>0$ is the strength of delayed feedback, $b>0$ is the decay rate, and $c$ is the shape parameter of the delayed feedback. The initial condition is defined as $x(t)=\phi(t)$ for $t\in[-\tau,0)$ and is assumed to be constant unless otherwise specified. The MG can exhibit multiple positive LEs, and the effective dimension of the dynamics increases as $\tau$ grows Doyne Farmer (1982).
@@
where $\sigma$ represents the noise intensity and $W_{t}$ denotes a Wiener process. The parameters are fixed at the standard values $b=0.1$ and $c=10$, while $a$ and $\sigma$ serve as control parameters. The delay time is chosen to be sufficiently larger than the response time Mensour and Longtin (1998) $1/b$ and is set to $\tau=90$. Numerical simulations are performed using the Euler–Maruyama method with a time step of $\Delta t=0.01$, unless otherwise stated.
@@ II.2 Bifurcation in the stochastic Mackey–Glass equation
We briefly summarize the phenomenology of the dynamics in the MG and SMG. In the bifurcation diagram of the MG (Fig. 1 (a)), the Poincaré section, $dx/dt=0$, of the attractor in Eq. (1) is plotted as a function of the delayed feedback strength $a$ (blue dots). In the bifurcation diagram of the SMG (Fig. 1 (b)), projections of the random pullback attractors are shown as black dots (see Appendices A and B for details on pullback attractors and bifurcation diagram). The first and second LEs of the MG and SMG as functions of $a$ at $\sigma=0.0$ and $\sigma=0.15$ are shown in Fig. 1 (c) (blue and black, respectively). For the MG, the origin is stable for $0<a<0.1$. A pitchfork bifurcation occurs at $a=0.1$, giving rise to two symmetric fixed points with respect to the origin. As $a$ increases, a Hopf bifurcation emerges at $a\simeq 0.125$, producing a limit cycle. A bifurcation to chaos follows at $a\simeq 0.138$, with high-dimensional chaos appearing around $a\simeq 0.144$. In the SMG, the zero-crossing point of the largest LE occurs at $a=a_{c}\simeq 0.175$, beyond which stochastic chaosChekroun, Simonnet, and Ghil (2011) is observed. At $a=a_{d}\simeq 0.195$, the second largest LE becomes positive, indicating the emergence of high-dimensional stochastic chaos.
In summary, although the MG becomes multi-stable via a pitchfork bifurcation, the SMG exhibits a single attractor with two quasi-stable states as two deterministic attractors are connected by external noise. As $a$ increases, the random LEs of the SMG grow monotonically, leading to stochastic chaos and later to high-dimensional stochastic chaos. Notably, these stochastic bifurcations occur at significantly larger $a$ values than in the deterministic system.
Figure 1: (a) Bifurcation diagram of the MG with $\sigma=0$. (b) Bifurcation diagram of random pullback attractors of the SMG with $\sigma=0.15$. (c) First and second largest LEs for $\sigma=0$ (blue points) and $\sigma=0.15$ (black points). In the bifurcation diagram, random pullback attractors are computed using $1.0\times 10^{2}$ constant initial functions uniformly distributed in [1,1][-1,1] with a pullback time $t_{p}=4\times 10^{4}$. We designate $a_{1}=0.14$ as a representative example of a weakly nonlinear regime and $a_{2}=0.30$ as that of a strongly nonlinear regime. At $a=a_{1},a_{2}$, low- and high-dimensional deterministic chaos are observed, respectively. The values, $a_{c}$ and $a_{d}$, indicate the zero-crossing points of the first and second largest LEs in the presence of noise. The LEs and the bifurcation diagram were computed using a time step of $\Delta t=0.1$ and a resolution defined as $N:=\tau/\Delta t=900$. Note that for delay differential equations, the largest LEs are known to be well approximated with a limited resolution of $N=\tau/\Delta t\sim 100$Wernecke, Sándor, and Gros (2019).
@@ II.3 Stochastic resonance in the stochastic Mackey–Glass equation
Beyond the deterministic pitchfork bifurcation point ($a>0.1$), SR appears as switching dynamics between two quasi-stable states, associated with point, periodic, and strange attractors in the MG. Figure 2 illustrates SR for a regime with deterministic periodic attractors. Without noise, two symmetric limit cycles exist with respect to the origin (Fig. 2 (a)), each with a period of approximately $2\tau$. In the presence of noise, switching between these limit cycles emerges, and periodicity is enhanced at an optimal noise intensity. Figure 2 (c) shows the power spectra with and without noise (black and blue lines, respectively). Under optimal noise intensity, the peaks at $f=f_{n}^{*}\simeq n/\tau$ $(n=1,2,\ldots)$ reveal clear harmonics. The power at the primary resonant frequency $f_{1}^{*}$ reaches a maximum at the optimal noise intensity $\sigma=\sigma^{*}$, confirming the presence of SR. We refer to $\sigma^{*}$ as the resonance point (Fig. 2 (d)). For $\tau$ sufficiently larger than $1/b$, SR is observed over a wide range of parameters, and the primary resonant frequency is approximately given by $1/\tau$.
Figure 2: (a) Two deterministic attractors projected onto $(x(t),x(t-\tau))$ plane for $a=0.126$. The inset shows a magnified view of the limit cycle. (b) Time series for the deterministic case, $\sigma=0$ (blue dots), and the resonance case, $\sigma=\sigma^{*}$ (black dots). (c) Power spectra for $\sigma=0$ (blue lines) and the optimal noise intensity $\sigma=\sigma^{*}$ (black lines). (d) Power at the primary resonant frequency $f_{1}^{*}$ as a function of $\sigma$.
@@ III.1 Stable and chaotic SR
To characterize the dynamical properties of SR, we compute the random LEs of the SMG (see Appendix A for the definition of random LE). We examine two representative cases: one where the deterministic attractors exhibit low-dimensional chaos ($a=a_{1}=0.14$), and one where they exhibit high-dimensional chaos ($a=a_{2}=0.30$). In both cases, switching dynamics between two quasi-stable states appears, and the power spectrum peaks at $f^{*}_{n}\simeq n/\tau$ $(n=1,2,\ldots)$ are enhanced at the resonance point (Fig. 3). The sign of the largest random LE at the resonance point distinguishes the two types of SR: a negative exponent corresponds to stable SR (Fig. 3 (b)), whereas a positive exponent corresponds to chaotic SR (Fig. 3 (d)).
Figure 3: Power spectra for $\sigma=0$ (blue lines) and $\sigma=\sigma^{*}$ (black lines) at (a) $a=a_{1}$ and (c) $a=a_{2}$. Power at the primary resonant frequency $f_{1}^{*}$ (black points) as a function of $\sigma$ at (b) $a=a_{1}$ and (d) $a=a_{2}$. The largest random LE at $a=a_{1}$ and the largest six random LEs at $a=a_{2}$ are also shown in (b) and (d), respectively (red points). The random LEs were computed using a time step of $\Delta t=0.1$ and a resolution defined as $N:=\tau/\Delta t=900$.
In the case of stable SR, increasing the noise intensity causes $\lambda_{1}$ to become negative at $\sigma=\sigma_{0}$, at which point the chaotic attractor collapses into a random point attractor. As $\sigma$ is increased further, $\lambda_{1}$ attains a local maximum near the resonance point. Because the zero-crossing point $\sigma_{0}$ precedes the resonance point $\sigma^{*}$, chaos vanishes at resonance, rendering the system effectively analogous to the classical SR model with a double-well potential and periodic forcing Benzi, Sutera, and Vulpiani (1981). For sufficiently large $\sigma$, $\lambda_{1}$ decreases again, as trajectories stay almost always in contracting regions outside the attractors. In the case of chaotic SR, $\lambda_{1}$ initially increases and peaks near the resonance point. Here, $\sigma^{*}$ precedes $\sigma_{0}$, indicating that SR occurs while the system still possesses a positive random LE and thus retains stochastic chaos. Chaotic SR can be regarded as a class of noise-induced phenomena that enhances the characteristic periods of stochastic chaos. The ordering of $\sigma^{*}$ and $\sigma_{0}$ provide a clear diagnostic: $\sigma_{*}>\sigma_{0}$ for stable SR and $\sigma_{*}<\sigma_{0}$ for chaotic SR.
@@ III.2 Visualization of stable and chaotic SR
Previous studies on SR in chaotic systems have generally not distinguished chaotic SR from stable SR Anishchenko, Neiman, and Safanova (1993); Anishchenko, Safonova, and Chua (1992); Carroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Krawiecki et al. (2001); Jüngling et al. (2008); Matyjaśkiewicz et al. (2003). In this work, we characterize stochastic dynamical structures, including stochastic chaos, within the framework of random dynamical systems. This distinction becomes evident when the dynamics are visualized using space–time representations Arecchi et al. (1992); Yanchuk and Giacomelli (2017) and random pullback attractors (see Appendix A for the construction of random pullback attractors).
A space–time representation maps the time series onto a two-dimensional plane by introducing a discrete step $n\in\mathbb{Z}$ and a memory space $s\in[0,\tau)$. Both stable SR and chaotic SR appear as traveling waves in the memory space $s$ (Fig. 4, left column). The angle $\theta$, defined as the propagation direction relative to the orthogonal direction, decreases with increasing delayed feedback strength $a$, and approaches zero as $\tau$ increases. Consequently, the angle $\theta_{1}$ for stable SR is larger than $\theta_{2}$ for chaotic SR. Notably, chaotic SR exhibits pronounced periodicity, with chaotic internal structures within the traveling waves, indicating the coexistence of resonance and chaos in stochastic dynamics.
When the largest random LE is negative, the random pullback attractor collapses to a random point attractor since initially close trajectories converge and eventually synchronize. Denoting the distribution of all possible trajectories at time $t$ by $\rho_{t}$, stable SR can be interpreted as pseudo-periodic motion of $\rho_{t}$ on the random point attractor (Fig. 4 (b)). Conversely, when the random LEs are positive, the random pullback attractor becomes a random strange attractor. In this case, chaotic SR corresponds to pseudo-periodic motion of $\rho_{t}$ on a random strange attractor (Fig. 4 (d)). Such pseudo-periodic motion of $\rho_{t}$ is related to the problem of statistical periodicity Lasota and Mackey (1987); Losson and Mackey (1995). The precise relationship between chaotic SR and statistical periodicity remains an open question and constitutes an interesting direction for future study.
@@
Figure 4: Space–time representation at the resonance point for (a) stable SR ($a=a_{1},\sigma=0.162$) and (c) chaotic SR ($a=a_{2},\sigma=0.104$). In both cases, traveling waves appear in the memory space $s\in[0,\tau)$. The angle $\theta$ denotes the propagation direction of the traveling wave relative to the orthogonal direction, defined by $\tan\theta=\epsilon$, where $\epsilon$ represents the deviation of the resonant period (see Eq. (4)). From the measured angles $\theta_{1}$ (stable SR) and $\theta_{2}$ (chaotic SR), we obtain $\theta_{2}/\theta_{1}\simeq 0.786$, indicating that wave propagation is slower in the chaotic SR regime. Panels (b) and (d) show snapshots of the random pullback attractor at the resonance point computed from $1.0\times 10^{5}$ constant initial functions uniformly distributed on [1,1][-1,1] with a pullback time $t_{p}=2\times 10^{4}$, projected onto $(x(t),x(t-\tau))$ plane.